# FoleyBench: A Benchmark For Video-to-Audio Models

### v1

@@ 1 Introduction
The task of video-to-audio generation (V2A) seeks to synthesize realistic, temporally aligned audio directly from visual input, sometimes conditioned on an additional text caption for finer control. A key downstream use case in the creative industry (e.g., film, games) is Foley—the creation of sound effects such as object impacts, or ambient noises that must be both semantically appropriate and temporally synchronized with visible events. Foley is distinct from speech and music, which are typically added in other stages of production. Foley specifically refers to non-speech, non-music sound effects (like footsteps or object impacts) that are causally linked to visual actions. Recent work has rapidly advanced V2A generation to support Foley scenarios, with a diverse family of methods being developed, including autoregressive, diffusion-based, flow-matching-based, and ControlNet-based models [Mei2023-FoleyGen, viertola2025temporally, luo2023diff, xu2024video, benita2025cafa, cheng2025mmaudio, multifoley2024].
@@ 3 Experimental Setup
We benchmark FoleyBench against a diverse set of recent V2A generation models, across a range of architectures. Our evaluation includes an autoregressive model (V-AURA [viertola2025temporally]), which predicts audio tokens sequentially; several diffusion-based approaches (DiffFoley [luo2023diff], VTA-LDM [xu2024video], FoleyCrafter [zhang2024foleycrafter], and LOVA [cheng2025lova]), which iteratively refine audio from noise and often emphasize temporal alignment or long-form synthesis; and adapter-based methods such as CAFA [benita2025cafa], which guide a pretrained text-to-audio generator through additional control modules. We also consider transformer-based masked prediction models (MaskVAT [pascual2024masked], SpecMaskFoley [zhong2025specmaskfoley], and Seeing&Hearing [xing2024seeing]), which adapt the MaskGIT framework to audio by filling in masked spectrogram regions and in some cases leverage test-time optimization for better grounding. We also have V2A-Mapper [wang2024v2a] that uses embedding translation, which projects CLIP video embeddings into the CLAP audio–text space to condition an AudioLDM backbone without extensive retraining. Finally, we evaluate two large-scale, jointly-trained models: MultiFoley [multifoley2024], a diffusion transformer trained on internet videos and high-quality SFX libraries, and MMAudio [cheng2025mmaudio], a large flow-matching model that jointly trains across video, text, and audio modalities. Together, these baselines reflect the breadth of strategies explored in V2A research—from lightweight adapters to fully joint multimodal models.
@@ Appendix A Baselines
MultiFoley. Chen et al. [multifoley2024] introduce a diffusion transformer-based model designed for multimodal control (text, audio, and video). It is trained on a combination of internet videos and professional, high-quality SFX libraries to produce full-bandwidth (48kHz) audio.

### v2

@@ 1 Introduction
The task of video-to-audio generation (V2A) seeks to synthesize realistic, temporally aligned audio directly from visual input, sometimes conditioned on an additional text caption for finer control. A key downstream use case in the creative industry (e.g., film, games) is Foley—the creation of sound effects such as object impacts, or ambient noises that must be both semantically appropriate and temporally synchronized with visible events. Foley is distinct from speech and music, which are typically added in other stages of production. Foley specifically refers to non-speech, non-music sound effects (like footsteps or object impacts) that are causally linked to visual actions. Recent work has rapidly advanced V2A generation to support Foley scenarios, with a diverse family of methods being developed, including autoregressive, diffusion-based, flow-matching-based, and ControlNet-based models [Mei2023-FoleyGen, viertola2025temporally, luo2023diff, xu2024video, benita2025cafa, cheng2025mmaudio, chen2025video].
@@ 3 Experimental Setup
We benchmark FoleyBench against a diverse set of recent V2A generation models, across a range of architectures. Our evaluation includes an autoregressive model (V-AURA [viertola2025temporally]), which predicts audio tokens sequentially; several diffusion-based approaches (DiffFoley [luo2023diff], VTA-LDM [xu2024video], FoleyCrafter [zhang2024foleycrafter], and LOVA [cheng2025lova]), which iteratively refine audio from noise and often emphasize temporal alignment or long-form synthesis; and adapter-based methods such as CAFA [benita2025cafa], which guide a pretrained text-to-audio generator through additional control modules. We also consider transformer-based masked prediction models (MaskVAT [pascual2024masked], SpecMaskFoley [zhong2025specmaskfoley], and Seeing&Hearing [xing2024seeing]), which adapt the MaskGIT framework to audio by filling in masked spectrogram regions and in some cases leverage test-time optimization for better grounding. We also have V2A-Mapper [wang2024v2a] that uses embedding translation, which projects CLIP video embeddings into the CLAP audio–text space to condition an AudioLDM backbone without extensive retraining. Finally, we evaluate two large-scale, jointly-trained models: MultiFoley [chen2025video], a diffusion transformer trained on internet videos and high-quality SFX libraries, and MMAudio [cheng2025mmaudio], a large flow-matching model that jointly trains across video, text, and audio modalities. Together, these baselines reflect the breadth of strategies explored in V2A research—from lightweight adapters to fully joint multimodal models.
@@ Appendix A Baselines
MultiFoley. Chen et al. [chen2025video] introduce a diffusion transformer-based model designed for multimodal control (text, audio, and video). It is trained on a combination of internet videos and professional, high-quality SFX libraries to produce full-bandwidth (48kHz) audio.